In an engine, an almost ideal gas is compressed adiabatically to half its volume. In doing so, 2900 J of work is done on the gas. How much heat flows into or out of the gas? Q- Submit Request Answer Part B What is the change in internal energy of the gas? Submit Request Answer
Answer #1

a)

Heat flows into or out of the gas is zero i.e

Q=0 J

since process is adiabatic

b)

The change in internal energy is

\Delta E_{th}=Q-W=0-(-2900)=2900 J